Use reduction of order to solve ty + 2ty - 2y yi = t is a solution of the differential equation. = 0 if it is known that
Answer #1

Solution: Given differential equation is

t^2y''+2ty'-2y=0.......(1)

and

Th = t

assume that a second solution will have the form

y_2(t)=v(t)y_1(t)

\Rightarrow y_2(t)=vt

\Rightarrow y_2'(t)=v+tv'

\Rightarrow y_2''(t)=v'+v'+tv''=2v'+tv''

putting these values in (1), we get

t^2(2v'+tv'')+2t(v+tv')-2vt=0

\Rightarrow 2t^2v'+t^3v''+2tv+2t^2v'-2vt=0

\Rightarrow t^3v''+4t^2v'=0

\Rightarrow tv''+4v'=0

consider,

v'=w\ \& \ v''=w'

we get

tw'+4w=0

\Rightarrow tw'=-4w

\Rightarrow \frac{w'}{w}=-\frac{4}{t}

By integrating, we get

log(w)=-4log(t)+log(c)

\Rightarrow log(w)=log(t^{-4})+log(c)

\Rightarrow log(w)=log\left ( \frac{c}{t^4} \right )

taking antilog

\Rightarrow w=\frac{c}{t^4}

Substituting back w=v' , we get

v'=\frac{c}{t^4}

integrate both sides

\Rightarrow v=-\frac{c}{3t^3}+d

By choosing c=-3 \ \ \& \ \ d=0 , we have

v=\frac{1}{t^3}

Thus,

y_2(t)=\frac{1}{t^3}t

\Rightarrow y_2(t)=\frac{1}{t^2}

Hence the general solution is given by

y(t)=c_1y_1(t)+c_2y_2(t)

\Rightarrow y(t)=c_1t+\frac{c_2}{t^2}

Which is the required solution.

This complete the solution.